Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows:

Standard
Hours
Standard Rate
per Hour
Standard
Cost
30 minutes $5.80 $2.90

During August, 10,350 hours of direct labor time were needed to make 19,100 units of the Jogging Mate. The direct labor cost totaled $57,960 for the month.

Required:

1. What is the standard labor-hours allowed (SH) to makes 19,100 Jogging Mates?

2. What is the standard labor cost allowed (SH × SR) to make 19,100 Jogging Mates?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

5. The budgeted variable manufacturing overhead rate is $4.30 per direct labor-hour. During August, the company incurred $49,680 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month.
Answer #1

Standard hours allowed = 19,100*0.5

= 9550

Standard cost allowed = 9550*5.80

= 55390

Labor spending variance = 55390-57960

= 2570 U

Labor Rate Variance = (SR-AR) *AH = (5.80-57960/10350)*10350

= 2070 F

Labor efficiency variance = (SH-AH) *SR = (19100*0.5 - 10350)*5.80

= 4640 U

Voh rate variance = (SR-AR) *AH = (4.30 - 49680/10350)*10350

= 5175 U

Voh efficiency variance = (SH-AH) *SR = (9550-10350)*4.30

= 3440 U